Prev Next

SW&IT Meeting Minutes 2019-05-13

Attendees: Andrew, David, Igor, Matthieu, Petr, Stephen, Xianyu, Yang

Outlook

Labor Reporting

Report auto-generated every, Do not edit this table, use entries in smartsheets.

20190506
LR20190506 - Bec, Matthieu 40.0
LR20190506 - Colson, Andrew 40.0
LR20190506 - Edmeades, David 24.0
LR20190506 - Sola, Igor 0.0
LR20190506 - Kubanek, Petr 0.0
LR20190506 - Hooper, Stephen 80.0
LR20190506 - Zhang, Yang 40.0
LR20190506 - Zhang, Xianyu 43.0

News from management

  • CH accepted our offer, starts July 5

Previous Action Items

Open topics

Team reports

MB

  • OC, no WPC, Management
  • node-lz4 fix
  • SDSS ctd,
  • dms ctd.
  • UITS colloc - looking good
  • vault r/w, ketiv support renewal talked to Buki
  • TMT executive s/w review, pretty good
next
  • dms ctd.
  • staff/system responsibilites

DE

  • Post-close-of-registration L4AO updates
  • LBTO site mockup work
  • UITS colocation proposal outline
  • New user startup for optical science collaborators
  • IssueTrak display tweaks

AC

  • AGW1 testing - IT7631
  • QFITSVIEWER (DPUSER -> LUCIMASKALIGN)
    • can't perform longslit alignments unless the user aligns more than one object along the slit
    • can't align on targets without a background image
    • centroid marker does not appear exactly where user clickes - rounds to the nearest pixel position
  • safety training
next:
  • Add radar image to http://info
  • NUC kickstart for CentOS 7 for FACSUM
  • Talk to science group about require ARGOS keywords

YZ

  • TCS stuff:
    • Finished a baseline version for SHARK authorization.
    • Updated the "irtc" program to include recent changes (ApplyWaffleAO command and new SHARK instruments).
    • Still need to fix some small details (mostly Qt message translation and display stuff).
    • Next to create github pull request for these changes.
    • Also do some more testing before concluding the development.
  • Working on Olga's LBC sky flats wiki page program.
  • AO-Cravat:
    • Haven't heard from Al regarding the AO onsky test data information. Once he reaches back with more info, will do some testing.
    • Once TCS developments are wrapped up, planning on continue to do some web interface development.

PK

next:

XZ

Last week
  • OVMS test on the telescope
  • AdSec SW changes for OVMS FLAO/SOUL
  • FLAO verification with new SwitchBUC FW and SW
  • Config flao-dxwfs for OVMS
  • Test flao-dxwfs for OVMS
This week

SH

  • Gene's new computer
  • Github repositories RPM
  • AGW1 in lab, SOUL network extension
  • shutdown SSH access to mountain archive
  • SCADA

next:
  • RPM review to see if we need SE policy sets
  • Fedora 30 burn
  • SCADA view
  • Bucki SSD

IS


Last week

a handful of Havasupai pics
  • LUCI
    • Preparing a fix on issuetrak #7642 (RTD aladin bug? on reading X/Y coordinates).
    • Also working on a fix to #7473, RTD and back end maybe to be tested on one of the engineering nights in May. I'd like to fix this without changing existing code and without adding extra load on the backend. It is harder than originally planned, might drop it and replace it with some other fix, such as a warning message.

This week
  • LUCI
    • I'll keep working on the LUCI fix for a couple more days and then try to have a release ready.
  • OT
    • Andrew C. and Michelle put together a spreadsheet with OT issuetraks. Have meet with them and go over it to scope it.
Topic revision: r10 - 13 May 2019, IgorSola
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback